OpenStudy (anonymous):

Between which two integers does square root of 178 lie? A.13&14 B. 17&18 C. 88 & 89 D. 177 & 179

OpenStudy (anonymous):

Ok, so there are many methods you could probably use but a easy one I just thought of is, if you take all of the following numbers you're given and multiply them together, and see what you get. And if the number you are looking for is between them, that's your answer. For example: 17 * 17 = 289, 18*18 = 324, that's not even close to being between 178. You get the idea?

OpenStudy (anonymous):

\[17*17 = 289 \implies \sqrt{289} = 17\]

OpenStudy (anonymous):

ohh okay i get it, Thanks!:)

OpenStudy (anonymous):

Cool :)